Earth and Artistry, Inc. provides commercial landscaping services. Sasha Cairns, the firm’s owner, wants to develop cost estimates that she can use to prepare bids on jobs. After analyzing the firm’s costs, Cairns has developed the following preliminary cost estimates for each 1,000 square feet of landscaping.

Direct material $ 460
Direct labor (5 direct-labor hours at $10 per hour) 50
Overhead (at $16 per direct-labor hour) 80
Total cost per 1,000 square feet $ 590

  

Cairns is quite certain about the estimates for direct material and direct labor. However, she is not as comfortable with the overhead estimate. The estimate for overhead is based on the overhead costs that were incurred during the past 12 months as presented in the following schedule. The estimate of $16 per direct-labor hour was determined by dividing the total overhead costs for the 12-month period ($640,000) by the total direct-labor hours (40,000).

Regular Direct- Overtime Direct- Total Direct-
Total Overhead Labor Hours Labor Hours* Labor Hours
January $ 49,000 2,810 190 3,000
February 47,000 3,270 40 3,310
March 54,000 3,500 60 3,560
April 48,000 3,000 250 3,250
May 60,000 3,470 460 3,930
June 58,000 3,440 770 4,210
July 60,000 3,400 620 4,020
August 48,000 2,670 310 2,980
September 49,000 2,930 40 2,970
October 56,000 2,750 30 2,780
November 59,000 3,160 20 3,180
December 52,000 2,590 220 2,810
Total $ 640,000 36,990 3,010 40,000

*The overtime premium is 50 percent of the direct-labor wage rate.

Cairns believes that overhead is affected by total monthly direct-labor hours. Cairns decided to perform a least-squares regression of overhead (OH) on total direct-labor hours (DLH). The following regression formula was obtained.

OH = $32,207 + $6.34 DLH

Required:

  1. Using the overhead formula that was derived from the least-squares regression, determine a total variable-cost estimate for each 1,000 square feet of landscaping.

  2. Cairns has been asked to submit a bid on a landscaping project for the city government consisting of 66,000 square feet. Cairns estimates that 50 percent of the direct-labor hours required for the project will be on overtime. Calculate the incremental costs that should be included in any bid that Cairns would submit on this project. Use the overhead formula derived from the least-squares regression.

  1. After attending a seminar on activity-based costing, Cairns decided to further analyze the company’s activities and costs. She discovered that a more accurate portrayal of the firm’s cost behavior could be achieved by dividing overhead into three separate pools, each with its own cost driver. Separate regression equations were estimated for each of the cost pools, with the following results.

OH1 = $10,700 + $3.50 DLH,

where DLH denotes direct-labor hours

OH2= $9,100 + $13.90 SFS,

where SFS denotes the number of square feet of turf seeded (in thousands)

OH3 = $8,400 + $7.00 PL,

where PL denotes the number of individual plantings (e.g., trees and shrubs)

Assume that 5 direct-labor hours will be needed to landscape each 1,000 square feet, regardless of the specific planting material used.

     a. Suppose the landscaping project for the city will involve seeding all 66,000 square feet of turf and planting 90 trees and shrubs. Calculate the incremental variable overhead cost that Cairns should include in the bid.

     b. Recompute the incremental variable overhead cost for the city’s landscaping project assuming half of the 66,000-square-foot landscaping area will be seeded and there will be 240 individual plantings. The plantings will cover the entire 66,000-square-foot area.

Complete this question by entering your answers in the tabs below.

1) Using the overhead formula that was derived from the least-squares regression, determine a total variable-cost estimate for each 1,000 square feet of landscaping. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Total variable-cost ????????????

2.Cairns has been asked to submit a bid on a landscaping project for the city government consisting of 66,000 square feet. Cairns estimates that 50 percent of the direct-labor hours required for the project will be on overtime. Calculate the incremental costs that should be included in any bid that Cairns would submit on this project. Use the overhead formula derived from the least-squares regression.  (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Show less

Minimum bid??????????????

3. Suppose the landscaping project for the city will involve seeding all 66,000 square feet of turf and planting 90 trees and shrubs. Calculate the incremental variable overhead cost that Cairns should include in the bid. (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Total incremental variable overhead ?????????????

4.Recompute the incremental variable overhead cost for the city’s landscaping project assuming half of the 66,000-square-foot landscaping area will be seeded and there will be 240 individual plantings. The plantings will cover the entire 66,000-square-foot area. (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Total incremental variable overhead ?????????????????????????

THE QUESTION IS COMPLETE IT DOESNT GIVE MORE INFORMATION!!!
Answer #1
1 Regression formula:
OH = $32207 + $6.34 DLH
Variable overhead cost=$ 6.34 per DLH
Total variable cost:
$
Direct material 460
Direct labor (5*10) 50
Variable overhead (5*6.34) 31.7
Total 541.7
2 Landscape in thousands of square feet=66000/1000=66
Incremental cost:
$
Direct material (66*460) 30360
Direct labor (66*50) 3300
Variable overhead (66*6.34) 418.44
Overtime (Note:1) 2475
Total 36553.44
Note:1
Overtime in hours=Total DLH*50%=66*5*50%=165 DLH
Overtime rate per hour=DL wage rate+(DL wage rate*50%)=10+(10*50%)=10+5=$ 15
Overtime cost=Overtime in hours*Overtime rate per hour=165*15=$ 2475
3 Incremental variable overhead cost:
$
OH1 (66*5*3.50) 1155
(Landscape in 1000 sq ft*DLH per 1000 sq ft*3.50)
OH2 (66*13.90) 917.4
(Landscape in 1000 sq ft*13.90)
OH3 (90*7) 630
(Number of individual plantings*7)
Total 2702.4
4 Number of square feet of turf seeded (in thousands)=(66000*50%)/1000=33
Incremental variable overhead cost:
$
OH1 (33*5*3.50) 577.5
(Landscape in 1000 sq ft*DLH per 1000 sq ft*3.50)
OH2 (33*13.90) 458.7
(Landscape in 1000 sq ft*13.90)
OH3 (240*7) 1680
(Number of individual plantings*7)
Total 2716.2